OpenStudy (deadshot):

What are the zeros of the polynomial function: f(x) = (x - 4)(2x + 1)(x + 7)?

OpenStudy (harsimran_hs4):

you have already done it yourself (x - 4)(2x + 1)(x + 7) = 0 for what value of x?

Parth (parthkohli):

Either \(x - 4 = 0 \quad \Rightarrow \quad x = 0 + 4 \). Or \(2x + 1 = 0 \quad \Rightarrow \quad \cdots\)

OpenStudy (harsimran_hs4):

put each factor separately equal to 0 and try finding the answer just as @ParthKohli has done

OpenStudy (deadshot):

\[4, -\frac{ 1 }{ 2 }, and -7\] right?

Parth (parthkohli):

Yes.

OpenStudy (deadshot):

Thanks for the help!

Parth (parthkohli):

No problem!